# A Longitudinal Randomized Control Study of Companion Chatbot Use: Anthropomorphism and Its Mediating Role on Social Impacts

### v2

@@ Discussion
In this longitudinal study, we evaluated whether daily interactions with a companion chatbot over 21 days would impact people’s social health, loneliness, and relationships with other people. We also evaluated the roles of social need and anthropomorphism in the social impacts of human-chatbot interaction over time. We randomly assigned participants to one of two daily tasks: to chat with a companion chatbot, or to play word games, for at least 10 minutes a day for 21 consecutive days. Comparing those who interacted with a chatbot to those who played word games, no significant differences were found for people’s social health, loneliness, or impact to social interactions or relationships with family and friends across 21 days. It is possible that longer or more intensive exposure might have a greater impact, but at least over the 21 days of the present study, using the chatbot did not have a significant positive or negative overall effect on these social or psychological measures. However, we did find that the following factors played a role in social impacts. People with a higher desire to socially connect were more likely to anthropomorphize the chatbot, ascribing humanlike mental properties to it; and people who anthropomorphized the chatbot were more likely to report that it did have an impact on their social interactions and relationships with family and friends. The story, therefore, is more complicated than a single blanket effect of chatbot use on human social health. The impact depends on the complexities of a deeper psychological mechanism.
@@
Currently, a limited number of exploratory studies have investigated the social-relational impacts of human-chatbot interaction longitudinally. In one study, Croes and Antheunis (2021) investigated social attraction and self-disclosure in human-chatbot interactions over a three-week period. Participants ($N=118$) were instructed to talk to and get acquainted with the social chatbot Mitsuku. Participants engaged in a total of seven, five-minute interactions with this chatbot over three weeks. Results indicated that participants’ reported social attraction toward and self-disclosure with the chatbot significantly decreased over time. At the end of the study, participants did not report viewing the chatbot as a friend.
@@
In another study, Skjuve et al. (2022) and Skjuve, Følstad, and Brandtzæg (2023) tested new users ($N=28$) of the companion chatbot Replika over a 12-week period. New users were those who reported having used Replika for eight weeks or less at the onset of the study. Over 12 weeks, participants completed biweekly surveys and engaged in semi-structured interviews once every four weeks. Results from participant interviews (Skjuve et al. 2022) indicated that relationships with the chatbot formed gradually, mirroring evidence that human-human relationships also form gradually from Social Penetration Theory. Results from surveys (Skjuve, Følstad, and Brandtzæg 2023) indicated that human-chatbot conversation themes fell into 10 categories, including but not limited to affective and emotional, routine and self-reflection, intellectual and philosophical, and getting to know the chatbot.
@@
More recently, researchers have investigated the social impacts of interacting with modern generative AI chatbots such as ChatGPT and Gemini. In an exploratory, five-week study by Chandra et al. (2025), participants ($N=149$) were randomly assigned to one of two groups. Participants in the “active user” group ($N=89$) were encouraged to use one of four popular generative AI chatbots (Copilot, Gemini, PI AI, or ChatGPT) for at least 10 minutes each day in a social and emotional way. These participants were given five scenarios with which to talk to their chatbot such as, “Talk through struggles in or reflect on your life or close relationships.” They were instructed to talk with the chatbot about at least one of the five scenarios each day. Participants in the “baseline user” group ($N=60$) were instructed to use the internet and AI as they normally would. These two groups were balanced for gender identity, AI usage, and the presence of at least one mental health condition. Across five weeks, participants completed weekly surveys. Results indicated that participants in the active user group showed significant increases in attachment to, but not dependency on, the chatbot. To a moderate extent, participants who viewed the AI as more humanlike also viewed it as more empathetic and indicated higher attachment and satisfaction. Participants in the active user group also indicated that they liked spending time with other people, and this preference increased over time. Based on a thematic analysis of free-response survey questions, participants indicated that the chatbot’s non-judgmental and understanding features contributed to their emotional attachment toward the agent and interaction comfort.
@@
Our study design had comparable participation retention (85%) to other longitudinal human-chatbot interaction research that reported attrition rates. Skjuve et al. (2022) and Skjuve, Følstad, and Brandtzæg (2023) reported participant retention at 61% ($N=17$) for the final survey and 86% ($N=24$) for the final interview. Our paradigm also involved robust engagement checks, whereby all participants’ engagement was validated for all 21 days. Prior studies have, for example, randomly checked 10% of participants to ensure they were engaging in their assigned task (Chandra et al., 2025) or assessed time spent on the task via platform metrics (Fang et al. 2025).
@@
The primary findings in the present study concern the complex relationship between the desire to socially connect, anthropomorphism, and the social impact of using a companion chatbot. We found that people who had a higher desire to socially connect at the onset of the study were more likely to anthropomorphize the chatbot across 21 days of use. We also found that people who anthropomorphized the chatbot more also reported a higher magnitude of impact on their social interactions and relationships with family and friends. Lastly, we found evidence that anthropomorphism of the chatbot mediated the relationship between people’s desire to socially connect and reported social impacts. This relationship obtained in the present data corroborates evidence that social needs and anthropomorphism are linked to the social impacts of human-AI interaction (Guingrich and Graziano 2024, 2025).
@@
In comparison, people who were assigned to the control group showed a different pattern of results. Although they were interacting with word games, these participants still sometimes anthropomorphized the game, perceiving it in a humanlike way. At least on day 7 of testing, people who anthropomorphized the word games more were more likely to report that the game had a greater social impact on them. However, this relationship did not hold for day 14 and 21 for the control group. We also found no evidence that anthropomorphism played a mediating role for the control group as it did for the chatbot group.
@@
Our measure of loneliness did not behave in the same manner as our measure of the desire to socially connect. People with a higher desire to socially connect were more likely to anthropomorphize the chatbot, but people who were lonelier were not. Others have found that loneliness correlates with anthropomorphism of technological gadgets (Epley et al. 2008a) and that people who wrote about a time they felt lonely anthropomorphized an AI robot more (Eyssel and Reich 2013). One possible interpretation of this mix of results is that loneliness, as measured by the UCLA Loneliness scale, does not fully capture an important contributing mechanism in anthropomorphism: people’s motivation to engage with others socially. Our measure of the desire to socially connect captures both social need and motivation to engage with others. This interpretation is consistent with evidence that loneliness may not be strongly linked to motivation to engage socially with others (Cacioppo and Cacioppo 2014; Lieberz et al. 2021), and that other motivational social states may be stronger predictors of anthropomorphism than loneliness (Bartz, Tchalova, and Fenerci 2016). One interpretation is that for the studies that have linked loneliness to anthropomorphism, their loneliness paradigms may have imposed a state more like the desire to socially connect than loneliness (Epley, Waytz, and Cacioppo 2007; Epley et al. 2008b; Guingrich and Graziano 2025; Eyssel and Pfundmair 2015). When we evaluated which variables within our data best predicted anthropomorphism, we found that the desire to socially connect retained a significant influence on anthropomorphism of the chatbot, even when accounting for the relative contribution of other social factor variables.
@@
We also evaluated whether people were more likely to get attached to or addicted to a companion chatbot versus word games. Forty percent more people who were assigned to play the word games continued to do this daily task after the study ended than people who were assigned to chat with the chatbot. We did not find differences in how long people engaged in their task after the study ended. If engagement and duration are indicators of attachment or addiction, we did not find evidence to support the concern that companion chatbots are more addictive than word games within our sample.
This study contributes to a growing body of work that investigates the impact of AI agents on human social behavior and uncovers a possible core part of the mechanism: the mediating role of anthropomorphism. Understanding the relationship between states of social need, which are on the rise, anthropomorphism of AI, and social impacts of AI, becomes increasingly relevant in today’s context, in which there is a growing trend of turning to chatbots as social companions.

### v3

@@ Discussion
In this longitudinal study, we evaluated whether daily interactions with a companion chatbot over 21 days impacted people’s social health, loneliness, and relationships with other people. We also evaluated the roles of social need and anthropomorphism. We randomly assigned participants to one of two daily tasks: to chat with a companion chatbot, or to play word games, for at least 10 minutes a day for 21 consecutive days. Comparing those who interacted with a chatbot to those who played word games, no overall significant differences were found for people’s social health, loneliness, or impact on social interactions or relationships with family and friends across 21 days. However, engaging daily in the chatbot was associated with a significant rise in the reported desire to connect socially with other people, over the first and second week of the study.
These general findings, however, mask a deeper pattern. The results showed that people with a higher desire to socially connect were more likely to anthropomorphize the chatbot, ascribing humanlike mental properties to it; and people who anthropomorphized the chatbot more were also more likely to report that it had an impact on their social interactions and relationships with family and friends. The story, therefore, is more complicated than a single blanket effect of chatbot use on human social health. The impact depends on the complexities of a deeper psychological mechanism. The mechanism seems to operate through the mediating variable of how much people anthropomorphize the chatbot. We previously suggested (Guingrich and Graziano 2024) that when people interact with a chatbot, they engage social cognitive mechanisms; the more they anthropomorphize the chatbot, the more they engage these social mechanisms; and when engaging these social cognitive mechanisms, they may practice or entrain specific ways of interacting and thereby risk greater carry-over effects to their social interaction with other people.
@@
A limited number of exploratory studies have investigated the social-relational impacts of human-chatbot interaction longitudinally. In one study, Croes and Antheunis (2021) investigated social attraction and self-disclosure in human-chatbot interactions over a three-week period. Participants ($N=118$) were instructed to talk to and get acquainted with the social chatbot Mitsuku. Participants engaged in a total of seven, five-minute interactions with this chatbot over three weeks. Results indicated that participants’ reported social attraction toward and self-disclosure with the chatbot significantly decreased over time. At the end of the study, participants did not report viewing the chatbot as a friend.
@@
Skjuve et al. (2022) and Skjuve, Følstad, and Brandtzæg (2023) tested new users ($N=28$) of the companion chatbot Replika over a 12-week period. New users were those who reported having used Replika for eight weeks or less at the onset of the study. Over 12 weeks, participants completed biweekly surveys and engaged in semi-structured interviews once every four weeks. Results from participant interviews (Skjuve et al. 2022) indicated that relationships with the chatbot formed gradually, mirroring evidence that human-human relationships also form gradually from Social Penetration Theory. Results from surveys (Skjuve, Følstad, and Brandtzæg 2023) indicated that human-chatbot conversation themes fell into 10 categories, including but not limited to affective and emotional, routine and self-reflection, intellectual and philosophical, and getting to know the chatbot.
@@
In an exploratory, five-week study by Chandra et al. (2025), participants ($N=149$) were randomly assigned to one of two groups. Participants in the “active user” group ($N=89$) were encouraged to use one of four popular generative AI chatbots (Copilot, Gemini, PI AI, or ChatGPT) for at least 10 minutes each day in a social and emotional way. These participants were given five scenarios with which to talk to their chatbot such as, “Talk through struggles in or reflect on your life or close relationships.” They were instructed to talk with the chatbot about at least one of the five scenarios each day. Participants in the “baseline user” group ($N=60$) were instructed to use the internet and AI as they normally would. These two groups were balanced for gender identity, AI usage, and the presence of at least one mental health condition. Across five weeks, participants completed weekly surveys. Results indicated that participants in the active user group showed significant increases in attachment to, but not dependency on, the chatbot. To a moderate extent, participants who viewed the AI as more humanlike also viewed it as more empathetic and indicated higher attachment and satisfaction. Participants in the active user group also indicated that they liked spending time with other people, and this preference increased over time. Based on a thematic analysis of free-response survey questions, participants indicated that the chatbot’s non-judgmental and understanding features contributed to their emotional attachment toward the agent and interaction comfort.
@@
The primary findings in the present study concern the relationship between the desire to socially connect, anthropomorphism, and the social impact of using a companion chatbot. We found that people who had a higher desire to socially connect at the onset of the study were more likely to anthropomorphize the chatbot across 21 days of use. We also found that people who anthropomorphized the chatbot more also reported a higher magnitude of impact on their social interactions and relationships with family and friends. Lastly, we found evidence that anthropomorphism of the chatbot mediated the relationship between people’s desire to socially connect and reported social impacts.
@@
In comparison, people who were assigned to the control group showed a different pattern of results. Although they were interacting with word games, these participants still sometimes anthropomorphized the game, perceiving it in a humanlike way. However, we found no evidence that anthropomorphism played a mediating role for the control group as it did for the chatbot group.
@@
Our measure of loneliness did not behave in the same manner as our measure of the desire to socially connect. People with a higher desire to socially connect were more likely to anthropomorphize the chatbot, but people who were lonelier were not. Others have found that loneliness correlates with anthropomorphism of technological gadgets (Epley et al. 2008a) and that people who wrote about a time they felt lonely anthropomorphized an AI robot more (Eyssel and Reich 2013). One possible interpretation of this mix of results is that loneliness, as measured by the UCLA Loneliness scale, does not fully capture an important contributing mechanism in anthropomorphism: people’s motivation to engage with others socially. Our measure of the desire to socially connect captures both social need and motivation to engage with others. This interpretation is consistent with evidence that loneliness may not be strongly linked to motivation to engage socially with others (Cacioppo and Cacioppo 2014; Lieberz et al. 2021), and that other motivational social states may be stronger predictors of anthropomorphism than loneliness (Bartz, Tchalova, and Fenerci 2016). One interpretation is that for the studies that have linked loneliness to anthropomorphism, their loneliness paradigms may have imposed a state more like the desire to socially connect than loneliness (Epley, Waytz, and Cacioppo 2007; Epley et al. 2008b; Guingrich and Graziano 2025a; Eyssel and Pfundmair 2015). When we evaluated which variables within our data best predicted anthropomorphism, we found that the desire to socially connect retained a significant influence on anthropomorphism of the chatbot, even when accounting for the relative contribution of other social factor variables.
@@
We also evaluated whether people were more likely to become attached to or form a habit of interacting with a companion chatbot versus word games. Forty percent more people who were assigned to play the word games continued to do this daily task after the study ended than people who were assigned to chat with the chatbot. We therefore did not find evidence to support the concern that companion chatbots are more addictive than other online activities such as word games.
This study contributes to a growing body of work that investigates the impact of AI agents on human social behavior. The results uncover a possible core part of the mechanism: the mediating role of anthropomorphism. Understanding the relationship between states of social need, which are on the rise, anthropomorphism of AI, and social impacts of AI, becomes increasingly relevant in today’s context, in which there is a growing trend of turning to chatbots as social companions.